OpenStudy (anonymous):

determine the domain for f(x)= x/(over) x^2-4

OpenStudy (campbell_st):

well have a look at the values of x you can't put into the denominator, as they will create a value of zero, which is undefined in the denominator so solve \[x^2 - 4 = 0\] these will be the values that are excluded from the domain of all real x

OpenStudy (anonymous):

does it equal two

OpenStudy (campbell_st):

there are two values, one of them is x = 2 whats the other...?

OpenStudy (anonymous):

the other is x=0?

OpenStudy (anonymous):

maybe, i'm not sure

OpenStudy (anonymous):

Factor

OpenStudy (anonymous):

x^2-4

OpenStudy (campbell_st):

nope can you factor \[x^2 - 4 = 0\] its the difference of 2 squares

OpenStudy (anonymous):

\[\Huge x^2-a^2=(x+a)(x-a)\]

OpenStudy (anonymous):

i don't understand, the first one is x=2 but the other one do i get by plugging in x^2-4 into that equation?

OpenStudy (campbell_st):

ok... what is the value of (-2)^2 =

OpenStudy (anonymous):

-4

OpenStudy (campbell_st):

well look at it this way what is \[-2 \times -2 = \]

OpenStudy (anonymous):

-4

OpenStudy (campbell_st):

multiplying 2 negatives gives a positive so the answer is 4

OpenStudy (anonymous):

oh oopsy sorry

OpenStudy (campbell_st):

so the solutions to the quadratic \[x^2 - 4 = 0\] is \[x = \pm 2\]

OpenStudy (anonymous):

Cause every number really has a positive and negative square root

OpenStudy (campbell_st):

so those values are excluded from the domain. so the domain of the function is all real x except -2 and 2

OpenStudy (anonymous):

oh, okay then. so its all the numbers except 2 and -2

OpenStudy (anonymous):

thank you